# Code Clone Detection via an AlphaFold-Inspired Framework

### v1

@@ Abstract.
Code clone detection, which aims to identify functionally equivalent code fragments, plays a critical role in software maintenance and vulnerability analysis. Substantial methods have been proposed to detect code clones, but they fall short in capturing code semantics or relying on language-specific analyzers. Inspired by the remarkable success of AlphaFold in predicting three-dimensional protein structures from protein sequences, in this paper, we leverage AlphaFold for code clone detection based on the insight that protein sequences and token sequences share a common linear sequential structure. In particular, we propose AlphaCC, which represents code fragments as token sequences to ensure multi-language applicability and adapts AlphaFold’s sequence-to-structure modeling capability to infer code semantics. The pipeline of AlphaCC goes through three steps. First, AlphaCC transforms each input code fragment into a token sequence and, motivated by AlphaFold’s use of multiple sequence alignment (MSA) to enhance contextual understanding, constructs an MSA from lexically similar token sequences. Second, AlphaCC adopts a modified attention-based encoder based on AlphaFold to model dependencies within and across token sequences. Finally, unlike AlphaFold’s protein structure prediction task, AlphaCC computes similarity scores between token sequences through a late interaction strategy and performs binary classification to determine code clone pairs. Comprehensive evaluations on three language-diverse datasets demonstrate AlphaCC’s applicability across multiple programming languages. On two semantic clone detection datasets, it consistently outperforms all baselines, showing strong semantic understanding. Moreover, AlphaCC maintains competitive efficiency, enabling practical usage in large-scale clone detection tasks.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Software Engineering; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06copyright: none
@@ 1. INTRODUCTION
Code clones, which refer to code fragments that are identical or equivalent in functionality, are frequently introduced through copy-and-paste practices or repeated implementation of similar program logic (Baker, 1995; Kim et al., 2005) in software engineering. While such reuse can reduce engineering effort and accelerate software development, it may also introduce hidden maintenance risks. Such risks materialize when duplicated code originates from buggy source code, as undetected defects propagate across the codebase, leading to accumulating technical debt and increasing maintenance cost (Kim et al., 2017). Furthermore, many buggy clones are not exact copies but modified versions of the original code, making the underlying bugs harder to detect and eliminate (Li and Ernst, 2012). To this end, code clone detection is essential for identifying and addressing issues that may exist in semantically equivalent code fragments distributed throughout the codebase.
@@
Various approaches have been proposed for code clone detection. Based on the intermediate representation used, these methods can be divided into four main categories (Jiang et al., 2007; White et al., 2016): text-based, token-based, Abstract Syntax Tree (AST)-based, and graph-based methods. The latter two categories of methods (AST‑based and graph‑based) extract structured intermediate representations (Jiang et al., 2007; Zhang et al., 2019; Xu et al., 2024; Yu et al., 2023; Zhao and Huang, 2018; Hua et al., 2020; Wu et al., 2020), such as abstract syntax trees and program dependency graphs, which are well-suited for detecting semantic clones (Wang et al., 2023), namely code fragments that perform the same or very similar functionality but differ syntactically (Yu et al., 2019). However, both methods rely on language-specific analyzers for structural extraction and suffer from low efficiency. The lack of corresponding analysis tools to support reliable construction of ASTs or program graphs is especially serious for some domain-specific languages (DSLs), such as SageMath (sag, 2025). Considering scenarios such as hardware design and data analytics, where the use of DSLs is steadily increasing (Damyanov and Sukalinska, 2015), AST-based and graph-based methods become impractical. In contrast, text-based and token-based methods employ relatively simple intermediate representations: the former treats source code as plain text, while the latter performs lexical analysis to transform source code into a sequence of tokens (Baker, 1997; Ankali and Parthiban, 2021; Feng et al., 2024). These lightweight representations offer advantages in efficiency and broad language applicability, but sacrifice the capability to capture code semantics (Sajnani et al., 2016; Kamiya et al., 2002). In summary, existing clone detection methods suffer from either limited capability for semantic clone detection or strong reliance on language-specific analyzers.
To address these limitations, we aim to develop a new code clone detection framework that features both semantic understanding capability and multi-language applicability. Our inspiration comes from AlphaFold (Jumper et al., 2021), which has achieved remarkable success in the biological domain by accurately predicting three-dimensional protein structures from protein sequences. The success of AlphaFold demonstrates its ability to abstract rich structural representations from linear symbolic sequences by capturing complex sequence dependencies that govern functional organization (Chen et al., 2024; Wang et al., 2024). Motivated by this, we explore whether AlphaFold can be applied to code clone detection, based on the fact that code fragments can likewise be represented as symbolic sequences, namely token sequences. Interestingly, token sequences and protein sequences share several key similarities:
  • Linear symbolic structure: Both are linear sequences of discrete symbols, i.e. tokens and amino acids, respectively.
  • Inner-sequence interaction: The meaning of each token or amino acid in a sequence depends on its sequence neighborhood.
  • Cross-sequence interaction: Homologous protein sequences are analogous to lexically similar token sequences, which provide context for learning underlying function.
Building on these similarities, we introduce a new framework to detect code clones that, on the one hand, benefits from a token-based representation without dependence on language-specific analyzers, and on the other hand, leverages the sequence-to-structure modeling capability of AlphaFold to capture code semantics.
@@
For an input protein sequence, AlphaFold first constructs a multiple sequence alignment (MSA) to incorporate information from homologous protein sequences, thereby enriching the input representation with co-evolutionary features (Rao et al., 2021). Then, an attention-based encoder, Evoformer, is employed to model complex dependencies, enabling information exchange both within individual sequences and across homologous sequences (Hu et al., 2022). The protein modeling process in AlphaFold inspires the design of a code clone detection framework that combines MSA construction and an attention-based encoder to enrich intermediate token representations and improve the capability to detect semantic clones.
@@
In this paper, we propose AlphaCC, an AlphaFold-inspired framework for code clone detection. Given a pair of code fragments, AlphaCC begins by preprocessing them into token sequences. It then proceeds through three main stages: a token semantic enhancer, an attention-based encoder, and a classification and optimization module. The token semantic enhancer aims to enrich token representation of code fragments. First, inspired by the MSA construction of AlphaFold, for each input token sequence, several lexically similar token sequences are retrieved from large-scale codebase together with their code context to construct a multiple sequence alignment of code (Code MSA). Then, AlphaCC introduces a novel mechanism that projects each token category into a type-specific semantic space, enabling the extraction of fine-grained, type-aware semantic features. Finally, an additional attention mechanism is applied to integrate information across these semantic spaces, enabling interactions between different token types. Building on these enriched token embeddings, an attention-based encoder adapted from AlphaFold’s Evoformer, termed Codeformer, is designed to integrate inner-sequence attention and cross-sequence attention, thereby capturing dependencies both within individual token sequence and across multiple token sequences. At the final stage, the classification and optimization module computes fine-grained similarity scores between code fragments via late interaction and performs binary classification to determine whether a given pair constitutes a code clone. To supervise this process, a margin-based loss function is employed as the optimization objective, increasing scores for semantically similar code fragment pairs while decreasing scores for dissimilar ones. It should be emphasized that the final stage is specific to the code clone detection task and is irrelevant to AlphaFold.
@@
To evaluate the effectiveness, multi-language applicability, and efficiency of AlphaCC, we conduct experiments on three widely used benchmarks for code clone detection: GCJ (gcj, 2017), BigCloneBench (big, 2020), and OJClone (Mou et al., 2016). Among them, GCJ and BigCloneBench are datasets based on Java, whereas OJClone consists of code written in C. The experimental results demonstrate that AlphaCC consistently outperforms all baseline methods across the three benchmarks, achieving clone detection accuracies of $97.4\%$, $96.0\%$, and $98.6\%$ on GCJ, BigCloneBench, and OJClone, respectively. Moreover, AlphaCC exhibits superior efficiency, achieving faster detection than all AST-based and graph-based approaches under identical computational settings.
@@
  • We observe several key similarities between protein sequences and token sequences, and creatively adapt the core idea of AlphaFold, namely sequence-to-structure modeling, to the task of code clone detection.
  • We propose AlphaCC, a code clone detection framework that adopts token representation and, inspired by AlphaFold, incorporates multiple sequence alignment construction and a modified attention-based encoder.
  • We conduct comprehensive experiments on three widely used benchmarks, demonstrating the effectiveness, multi-language applicability, and efficiency of AlphaCC.
The remainder of the paper is organized as follows. Section 2 reviews related work on code clone detection. Section 3 introduces the biological inspiration behind our method. Section 4 presents AlphaCC, an AlphaFold-inspired framework for code clone detection. Section 5 details the experimental setup and results to evaluate the effectiveness, multi-language applicability, and efficiency of AlphaCC. Section 6 discusses the possible interpretations of AlphaCC and examines potential threats to validity. Section 7 concludes this paper.
@@
2. RELATED WORK
This section introduces related works on code clone detection. According to different intermediate representations of code, existing code clone detection methods can be divided into four categories: text-based, token-based, AST-based, and graph-based.
Text-based methods treat source code as plain text and employ string matching techniques to identify clones. Johnson (Johnson, 1993) designs a prototype system based on the fingerprints of code to identify code clones. Baker proposes Dup (Baker, 1995), which detects duplicate code by converting code strings into parameterized suffix trees and then identifying matches within these trees. Despite their rapid detection speed, the absence of program syntax weakens text-based methods in handling complex syntactic or semantic clones.
AST-based methods use the abstract syntax trees of code snippets to detect code clones. Baxter et al. (Baxter et al., 1998) split the original AST into subtrees and compare the hash code of subtrees to detect code clones. Jiang et al. (Jiang et al., 2007) design DECKARD to compute characteristic vectors for AST subtrees and employ locality-sensitive hashing to group similar vectors, thereby identifying clones. Furthermore, some methods, such as CDLH (Wei and Li, 2017), ASTNN (Zhang et al., 2019), and DSFM(Xu et al., 2024), use deep learning models to encode AST into vectors and then calculate the similarity between vectors to detect code clones. While AST-based approaches can identify semantic code clones, they are often hindered by efficiency issues and limited multi-language applicability because of heavy reliance on language-specific AST analyzers.
For graph-based methods, they first analyze different graph representations, such as control flow graph (CFG) and program dependency graph (PDG), for code, and then further process the representation data to detect code clones. Krinke (Krinke, 2001) finds similar subgraphs in PDG to detect code clones. Wu et al. (Wu et al., 2020) design SCDetector, which uses centrality analysis to detect code clones from CFG. Similar to AST-based methods, graph-based approaches are effective in capturing semantic similarities. However, graph-based approaches face low detection efficiency, and their dependence on language-specific graph construction tools further restricts widespread adoption.
In summary, existing clone detection methods are constrained by insufficient ability of semantic clone detection or strong dependence on language-specific analyzers. These limitations motivate the design of our new framework, AlphaCC.
@@ 2. RELATED WORK
3. Biological Inspiration: From Protein Structure Prediction to Code Clone Detection
This section introduces the biological inspiration behind our approach. Section 3.1 draws an analogy between protein sequences and token sequences, while Section 3.2 compares the framework of AlphaFold and AlphaCC.
3.1. From Protein to Code: Anology of Sequence
We establish an analogy between protein sequences and token sequences from three key perspectives.
@@ 3.1. From Protein to Code: Anology of Sequence
First, as shown in Figure 1, protein sequences and token sequences share a common structural pattern: a linear symbolic structure. That is, both protein sequences and token sequences are linear sequences consisting of a finite number of discrete symbols—amino acids in the former and code tokens in the latter. Moreover, there are $20$ standard types of amino acids in protein sequences, such as Alanine, Cysteine, Lysine, and Proline. Similarly, for a given programming language, code tokens fall into a finite number of categories, such as modifiers, identifiers, keywords, and operators. Each token belongs to exactly one category, as determined by the language’s lexical analyzer.

### v2

@@ Abstract.
Code clone detection, which aims to identify functionally equivalent code fragments, plays a critical role in software maintenance and vulnerability analysis. Substantial methods have been proposed to detect code clones. However, they struggle to extract high-level program semantics directly from a single linear token sequence, leading to unsatisfactory detection performance. A similar single-sequence challenge has been successfully addressed in protein structure prediction by AlphaFold through learning rich representations from single-sequence inputs. Motivated by the successful resolution of the shared single-sequence challenge by AlphaFold, as well as the sequential similarities between proteins and code, we leverage AlphaFold for code clone detection. In particular, we propose AlphaCC, which represents code fragments as token sequences and adapts AlphaFold’s sequence-to-structure modeling capability to infer code semantics. The pipeline of AlphaCC goes through three steps. First, AlphaCC transforms each input code fragment into a token sequence and, motivated by AlphaFold’s use of multiple sequence alignment (MSA), novelly uses a retrieval-augmentation strategy to construct an MSA from lexically similar token sequences. Second, AlphaCC adopts a modified attention-based encoder based on AlphaFold to model dependencies within and across token sequences. Finally, unlike AlphaFold’s protein structure prediction task, AlphaCC computes similarity scores between token sequences through a late interaction strategy and performs binary classification to determine code clone pairs. Comprehensive evaluations on three datasets, particularly two semantic clone detection datasets, show that AlphaCC consistently outperforms all baselines, demonstrating strong semantic understanding. AlphaCC further achieves strong performance on instances where tool-dependent methods fail, highlighting its tool-independence. Moreover, AlphaCC maintains competitive efficiency, enabling practical usage in large-scale clone detection tasks.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Code clones, which refer to code fragments that are identical or equivalent in functionality, are frequently introduced through copy-and-paste practices or repeated implementation of similar program logic (Baker, 1995; Kim et al., 2005) in software engineering. While such reuse can reduce engineering effort and accelerate software development, it may also introduce hidden maintenance risks. Such risks materialize when duplicated code originates from buggy source code, as undetected defects propagate across the codebase, leading to accumulating technical debt and increasing maintenance cost (Kim et al., 2017; Shan et al., 2023; Xu et al., 2024). Furthermore, many buggy clones are not exact copies but modified versions of the original code, making the underlying bugs harder to detect and eliminate (Li and Ernst, 2012; Ebrahimi et al., 2019). For example, the OpenSSL Heartbleed vulnerability (CVE-2014-0160) (8) affected a wide range of systems due to direct reuse or partial cloning of the OpenSSL library (Walden, 2020). To this end, code clone detection is essential for identifying and addressing issues that may exist in semantically equivalent code fragments distributed throughout the codebase.
@@
Various approaches have been proposed for code clone detection. From the perspective of code representation, existing methods can be divided into two main categories: token-based and intermediate-representation (IR)-based (Hua et al., 2020; Feng et al., 2024). IR-based methods transform source code into structured intermediate representations, which typically take the form of Abstract Syntax Trees (ASTs) or program graphs. Existing approaches (Jiang et al., 2007; Zhang et al., 2019; Xu et al., 2024; Yu et al., 2023; Zhao and Huang, 2018; Hua et al., 2020; Wu et al., 2020) leverage such representations to effectively detect semantic clones (Wang et al., 2023), namely code fragments that perform the same or very similar functionality but differ syntactically (Yu et al., 2019). However, IR-based methods rely on third-party analyzers for structural extraction and suffer from low efficiency. In practice, limitations in the maturity and maintenance of third-party analyzers—such as insufficient support for certain programming language versions—can prevent IR-based methods from operating (see Section 5.5). Considering legacy software using older language versions, IR-based methods exhibit limited generalization and robustness. In contrast, token-based methods operate directly on the lexical representation of source code, where programs are treated as token sequences (Baker, 1997; Ankali and Parthiban, 2021; Feng et al., 2024). Without complex structural extraction pipelines, token-based methods are generally more efficient and tool-independent; however, they sacrifice the ability to capture code semantics (Sajnani et al., 2016; Kamiya et al., 2002). In summary, existing clone detection methods suffer from either limited capability for semantic clone detection or strong reliance on third-party analyzers. These limitations stem from a fundamental challenge in code clone detection: extracting high-level program semantics directly from a single linear token sequence.
A closely related challenge has long been studied in protein structure prediction: extracting three-dimensional structural information directly from a single linear protein sequence. In both domains, the shared single-sequence challenge arises: learning high-level representations from a single linear sequence. Recent advances in protein structure prediction—most notably AlphaFold (Jumper et al., 2021)—demonstrate that this shared challenge can be effectively overcome. By learning rich high-level representations from a single protein sequence, AlphaFold provides a successful solution to this long-standing limitation (Chen et al., 2024; Wang et al., 2024).
Beyond the shared challenge, we further observe sequential similarities between proteins and code from a representational perspective. Both can be represented as linear symbolic sequences—amino acid sequences and token sequences, respectively. Each sequence is composed of a limited set of symbols, for example, the 20 standard amino acids in proteins or the finite set of programming tokens in code.
Motivated by AlphaFold’s success in addressing the shared single-sequence challenge, together with the sequential similarities between proteins and code, we introduce a new framework to address the single-sequence challenge in code clone detection. The proposed framework adopts a token-based representation without dependence on third-party analyzers (that often get obsolete over time), and leverages the sequence-to-structure modeling capability of AlphaFold to capture code semantics.
@@
For an input protein sequence, since the limited information it contains is insufficient to reliably infer higher-level structural properties, AlphaFold first constructs a multiple sequence alignment (MSA). By retrieving homologous protein sequences, MSA enriches the input representation with co-evolutionary information (Rao et al., 2021). Then, in order to model complex sequence dependencies, an attention-based encoder, Evoformer, is employed, enabling information exchange both within individual sequences and across homologous sequences (Hu et al., 2022). The subsequent decoder module in AlphaFold, which maps learned representations to three-dimensional coordinates, is specific to protein structure prediction and is therefore not considered here. The protein modeling process in AlphaFold inspires the design of a code clone detection framework that combines MSA construction and an attention-based encoder to enrich token representations and improve the capability to detect semantic clones.
@@
In this paper, we propose AlphaCC, an AlphaFold-inspired framework for code clone detection. AlphaCC proceeds through three main stages: a token semantic enhancer, an attention-based Codeformer encoder, and a classification and optimization module. The token semantic enhancer aims to enrich token representation of code fragments. First, inspired by the MSA construction of AlphaFold, for each input token sequence, several lexically similar token sequences are retrieved from large-scale codebase together with their code context to construct a multiple sequence alignment of code (Code MSA). To the best of our knowledge, this is the first time a retrieval-augmented strategy is applied to code clone detection. Then, AlphaCC introduces a novel mechanism that projects each token category into a type-specific semantic space, enabling the extraction of fine-grained, type-aware semantic features. Finally, an additional attention mechanism is applied to integrate information across these semantic spaces, enabling interactions between different token types. Building on these enriched token embeddings, an attention-based encoder adapted from AlphaFold’s Evoformer, termed Codeformer, is designed to integrate inner-sequence attention and cross-sequence attention, thereby capturing dependencies both within individual token sequences and across multiple token sequences. At the final stage, the classification and optimization module computes fine-grained similarity scores between code fragments via late interaction and performs binary classification to determine whether a given pair constitutes a clone pair. To supervise this process, a margin-based loss function is employed as the optimization objective, increasing scores for semantically similar code fragment pairs while decreasing scores for dissimilar ones. It should be emphasized that the final stage is specific to the code clone detection task and is irrelevant to AlphaFold.
@@
To evaluate the effectiveness, tool-independence, and efficiency of AlphaCC, we conduct experiments on three widely used code clone detection benchmarks: GCJ (14), BigCloneBench (5), and OJClone (Mou et al., 2016), among which GCJ and OJClone focus on semantic clone. Experimental results show that AlphaCC consistently outperforms all baseline methods across the three benchmarks, achieving clone detection accuracies of $97.4\%$, $96.0\%$, and $98.6\%$ on GCJ, BigCloneBench, and OJClone, respectively. Notably, AlphaCC is, to the best of our knowledge, the first token-based approach reported to surpass IR-based methods, showing strong semantic understanding. Moreover, AlphaCC maintains strong performance on instances where tool-dependent methods fail to operate, highlighting its tool-independence. It also exhibits superior efficiency, achieving faster detection than all IR-based approaches under identical computational settings.
@@
  • We observe similarities between protein sequences and token sequences, and creatively adapt the core idea of AlphaFold, namely sequence-to-structure modeling, to the task of code clone detection.
  • We propose AlphaCC, a token-based code clone detection framework that adopts token representation and, inspired by AlphaFold, incorporates multiple sequence alignment construction and a modified attention-based encoder.
  • We conduct comprehensive experiments on three widely used benchmarks, demonstrating the effectiveness, tool-independence, and efficiency of AlphaCC.
The remainder of the paper is organized as follows. Section 2 reviews related work on code clone detection. Section 3 introduces the biological inspiration behind our method. Section 4 presents AlphaCC, an AlphaFold-inspired framework for code clone detection. Section 5 details the experimental setup and results to evaluate the effectiveness, tool-independence, and efficiency of AlphaCC. Section 6 discusses the possible interpretations of AlphaCC and examines potential threats to validity. Section 7 concludes this paper.
2. RELATED WORK
This section introduces related works on code clone detection. According to different intermediate code representations, existing code clone detection methods can be divided into two categories: token-based and IR-based.
IR-based methods analyze programs through structured intermediate representations that reflect program structure and semantics. According to the type of intermediate representations they rely on, IR-based methods can be further categorized into AST-based and graph-based methods.
AST-based methods use the abstract syntax trees of code fragments to detect code clones. Baxter et al. (Baxter et al., 1998) split the original AST into subtrees and compare the hash code of subtrees to detect code clones. Jiang et al. (Jiang et al., 2007) design DECKARD to compute characteristic vectors for AST subtrees and employ locality-sensitive hashing to group similar vectors, thereby identifying clones. Furthermore, some methods, such as CDLH (Wei and Li, 2017), ASTNN (Zhang et al., 2019), and DSFM(Xu et al., 2024), use deep learning models to encode AST into vectors and then calculate the similarity between vectors to detect code clones. While AST-based approaches can identify semantic code clones, they are often hindered by efficiency issues and limited tool-independence because of reliance on third-party AST analyzers.
For graph-based methods, they first analyze different graph representations, such as control flow graph (CFG) and program dependency graph (PDG), for code, and then further process the representation data to detect code clones. Krinke (Krinke, 2001) finds similar subgraphs in PDG to detect code clones. Wu et al. (Wu et al., 2020) design SCDetector, which uses centrality analysis to detect code clones from CFG. Similar to AST-based methods, graph-based approaches are effective in capturing semantic clones. However, graph-based approaches face low detection efficiency, and their dependence on third-party graph construction tools further restricts widespread adoption.
In summary, existing clone detection methods are constrained by insufficient ability of semantic clone detection or strong dependence on third-party analyzers. These limitations motivate the design of our framework, AlphaCC.
@@ 2. RELATED WORK
3. Biological Inspiration: From AlphaFold to AlphaCC
This section presents the shared design rationale underlying AlphaFold and AlphaCC, and draws an analogy between the frameworks of AlphaFold and AlphaCC to emphasize their connection and distinction.
AlphaFold is a representative end-to-end framework that takes a protein sequence as input and predicts its three-dimensional structure. As illustrated in Figure 2, the overall framework consists of three stages: MSA construction, the Evoformer module, and the structure module. Since the structure module in AlphaFold serves as a decoder that maps learned features to three-dimensional coordinates—a stage irrelevant to code clone detection—AlphaCC focuses on adapting the first two stages to extract semantic features from token sequences.
Accordingly, this section concentrates on the first two stages of these two frameworks. Section 3.1 discusses the adaptation of MSA construction, while Section 3.2 describes the adaptation of Evoformer Encoder.
3.1. MSA Construction Adaptation
@@
The adaptation of MSA Construction is motivated by two similarities between protein structure prediction and code clone detection.
@@
First, there exists a sequential similarity between protein sequences and token sequences. As shown in Figure 1, both sequences share a common structural pattern: a linear symbolic structure. That is, both protein sequences and token sequences are linear sequences consisting of finite sets of discrete symbols—amino acids in the former and code tokens in the latter. Moreover, each symbol belongs to a predefined category. Protein sequences consist of $20$ standard amino acid types, such as Alanine, Cysteine, and Proline. Similarly, for a given programming language, code tokens fall into a finite number of categories, such as identifiers, keywords, and operators. Each token belongs to exactly one category, as determined by the language’s lexical analyzer. Second, both protein structure prediction and code clone detection face an inherent challenge: a single sequence contains only limited information, which is insufficient to infer deeper properties—three-dimensional structure for proteins (Shih et al., 2012) and program functionality for code.
To address this limitation, AlphaFold introduces MSA to aggregate homologous protein sequences into an aligned representation, enabling the capture of evolutionary and structural patterns that cannot be inferred from a single sequence alone (Lupo et al., 2022; Rao et al., 2021; Zhang et al., 2024). Motivated by these similarities and the role of MSA in AlphaFold, AlphaCC introduces Code MSA to enrich the semantic representation of individual token sequences. For each target sequence, AlphaCC retrieves functionally similar code instances from a large-scale codebase and aligns them to form a Code MSA, providing a broader context than an isolated code fragment for subsequent representation learning. To further enrich token representation, AlphaCC additionally introduces type-specific projections and an interaction mechanism. Overall, Code MSA serves as the foundation for enhancing semantic understanding in AlphaCC.
3.2. Encoder Adaptation
The adaptation of Evoformer encoder stems from a shared phenomenon observed in protein sequences and token sequences: both sequences exhibit element interaction, where the meaning of each element is context-dependent and determined by its neighboring elements.